OpenStudy (anonymous):

Find the question of the line with slope -3/4 and passing through (-3,3) . help please ! I'm stuck !

OpenStudy (aripotta):

find the question? do you mean equation?

OpenStudy (anonymous):

yes, equation, i'm sorry

OpenStudy (aripotta):

what form does it have to be in?

OpenStudy (anonymous):

y=mx+b

OpenStudy (aripotta):

alright, let's put it in point-slope first. y - y1 = m(x - x1)

OpenStudy (aripotta):

so x1 is -3, y1 is 3, and m is -3/4

OpenStudy (aripotta):

y - 3 = -3/4(x + 3)

OpenStudy (aripotta):

can you put it in slope-intercept now?

OpenStudy (anonymous):

Thats where i was stuck :O

OpenStudy (aripotta):

ok, all you have to do is first distribute the -3/4 into the parentheses and then you can move stuff around

OpenStudy (aripotta):